Find the indicated limit, if it exists. (2 points)
limit of f of x as x approaches 5 where f of x equals 5 minus x when x is less than 5, 8 when x equals 5, and x plus 3 when x is greater than 5

0

8

3

The limit does not exist.

Answer by greenestamps(13203)



The limit as x approaches 5 from the left (x<5) is 5-x = 5-5 = 0; the limit as x approaches 5 from the left (x>=5) is 8.

By definition, since the limits from the two directions are not the same, the limit does not exist.